### v2

# Relational Quantum Mechanics: Ozawa’s Intersubjectivity Theorem as justification of the postulate on internally consistent descriptions

Abstract
The Ozawa’s Intersubjectivity Theorem (OIT) proved within quantum measurement theory supports the new postulate of relational quantum mechanics (RQM), the postulate on internally consistent descriptions. But from OIT viewpoint postulate’s formulation should be completed by the assumption of probability reproducibility. We remark that this postulate was proposed only recently to resolve the problem of intersubjectivity of information in RQM. In contrast to RQM for which OIT is a supporting theoretical statement, QBism is challenged by OIT.
keywords: Relational Quantum Mechanics, Ozawa Intersubjectivity Theorem, postulate on internally consistent descriptions, quantum instruments, measurement process, reproducibility of probability distribution, unitary evolution
1 Introduction
Recently Ozawa’s Intersubjectivity Theorem (OIT) [1] proved within quantum measurement theory (e.g., [2]-[8]) was used as objection to QBism [9], to its basic interpretational statement about the individual agent perspective on measurement’s outcomes [10]-[12]. (This privacy component of QBism was also criticized in my previous articles [13, 14], mainly from the probabilistic perspective.) Now QIT provides the strong argument that measurement’s outcome is intersubejctive and, although one can still interpret it as a private experience, such interpretation loses its elegance. QBists replied to my objection with two papers [15, 16] (the second one presents more elaborated counterargument).
In a few words the QBism’s counterargument is that compatible measurement devices $M_{1}$ and $M_{2}$ considered in OIT should be related not to two distinct observers $Q_{1}$ and $O_{2}$ as was done in my article [9], but to one observer $O$ who performs these two measurements. As was clearly explained in papers [15, 16], this interpretation of such measurements is consistent with the general ideology of QBism. The important impact of article [9] to QBism is that this article attracted the attention (at least some) of QBists to quantum measurement theory [2]-[8]. QBism intensively explores one of the basic elements of this theory - POVMs. But QBists didn’t employ the important part of the quantum measurement theory, namely, that every measurement described by completely positive instrument has a parallel process described by an interaction between the object and the apparatus, or more mathematically every abstract measuring process determines its completely positive instruments and that every completely positive instruments arises in this way. (This was proven by Ozawa in article [3].)
The discussions on the intersubjectivity problem in QBism stimulated my interest to this problem for other information interpretations of QM. A reviewer of one of my papers pointed that relational quantum mechanics (RQM) [17]- [21] was suffering of the same problem. It seems that it cannot be resolved within “old RQM”. Its resolution required invention of an additional postulate on internally consistent descriptions (postulate 6 in [21]). By treating this postulate within quantum instrument theory, one immediately see that OIT supports this postulate. So, in contrast to QBism which was challenged by OIT, RQM’s postulate 6 is mathematically justified via OIT.
However, the OIT view on postulate 6 leads to the conclusion that
  • on the one hand, to be supported by the quantum theory, this postulate should be constrained by the probability reproducibility condition,
  • on the other hand, it can be formulated in a stronger form.
As well as QBism, RQM can earn (for the moment, it is not clear how much?) from its framing withing the mathematical formalism of quantum measurement theory [2]-[8]. Besides the present paper, an attempt to proceed in this direction was done in article [22]. However, the conclusions of this article are debatable. I am not an expert in RQM and I am not able to evaluate whether the authors of [22] correctly understood the conceptual premises of RQM.
We start with a brief recollection of the basics of RQM and the problem of intersubjectivity in its framework. Here we follow article [21].
2 Relational quantum mechanics and the problem of intersubjectivity
As was stated by Rovelli [17], RQM is based on the idea that “in quantum mechanics different observers may give different accounts of the same sequence of events.” It is crucial that “RQM is built on strong naturalistic intuitions; therefore, in RQM, the term “observer” is understood in a broad sense, which allows that any physical system can be an “observer,” so we do not have to accept that consciousness plays any fundamental role.” [21]. In the last aspect RQM differs crucially from QBism. As is emphasized in article [21], RQM has many attractive features and it resolves the basic quantum paradoxes (but QBism do this as well).
“However, some problems remain; in particular, there is a tension between RQM’s naturalistic emphasis on the physicality of information and the inaccessibility of certain sorts of information in current formulations of RQM. Thus, in this article, we propose a new postulate for RQM which ensures that all of the information possessed by a certain observer is stored in physical variables of that observer and thus is accessible by measurement to other observers. The postulate of cross-perspective links makes it possible for observers to reach intersubjective agreement about quantum events that have occurred in the past, thus shoring up the status of RQM as a form of scientific realism and allowing that empirical confirmation is possible in RQM.
Adding this postulate requires us to update some features of the ontology of RQM, because it entails that not everything in RQM is relational.”[21]
For readers convenience, here the list of all RQM’s postulates is presented [21]:
  1. 1.
    Relative facts: Events, or facts, can happen relative to any physical system.
  2. 2.
    No hidden variables: Unitary quantum mechanics is complete.
  3. 3.
    Relations are intrinsic: The relation between any two systems $A$ and $B$ is independent of anything that happens outside these systems’ perspectives.
  4. 4.
    Relativity of comparisons: It is meaningless to compare the accounts relative to any two systems except by invoking a third system relative to which the comparison is made.
  5. 5.
    Measurement: An interaction between two systems results in a correlation within the interactions between these two systems and a third one; that is, with respect to a third system $W,$ the interaction between the two systems $S$ and $F$ is described by a unitary evolution that potentially entangles the quantum states of $S$ and $F.$
  6. 6.
    Internally consistent descriptions: In a scenario where $O_{1}$ measures $S,$ and $O_{2}$ also measures $S$ in the same basis, and $O_{2}$ then interacts with $O_{1}$ to “check the reading” of a pointer variable (i.e., by measuring $O_{1}$ in the appropriate “pointer basis”), the two values found are in agreement.
OIT matches perfectly postulate 6, justifies it mathematically on the basis of quantum instrument theory. Generally my picture is that RQM can earn a lot from merging with quantum instrument theory.
As was cited above in “RQM the term “observer” is understood in a broad sense, which allows that any physical system can be an “observer…”
In quantum instrument theory, “apparatus” plays the key role. I would suggest to extend its meaning and operate in this theory with the notion observer. The latter has the RQM meaning, i.e., any physical system can be an observer. I understand that such a step may induce critique from both sides, RQM and quantum instrument theory.
3 Ozawa Intresubjectivity Theorem
Von Neumann [2] described observables mathematically by Hermitian operators in Hilbert space ${\cal H}.$ Consider operators with totally discrete spectra $X\subset\mathbb{R}:A=\sum_{x\in X}xE_{A}(x),$ where $E_{A}(x)$ is projection on the eigensubspace for the eigenvalue $x.$ The Born rule determines the probabilities:
$P(A=x|\psi)=\langle\psi|E_{A}(x)|\psi\rangle.$
The indirect measurement scheme involves the following components
  • the states spaces ${\cal H}$ and ${\cal K}$ of the systems ${\cal S}$ and the apparatus ${\cal M}$ for measurement of some observable $A;$
  • the evolution operator $U=U(t)$ representing the interaction-dynamics for the system ${\cal S}+{\cal M};$
  • the meter observable $M$ giving outputs of the pointer of the apparatus ${\cal M}.$
Here the quantum observables $A$ and $M$ can be represented as PVMs, $E_{A}=(E_{A}(x)),E_{M}=(E_{M}(x)),$ where $E_{A}(x),E_{M}(x)$ are projections in Hilbert spaces ${\cal H}$ and ${\cal K}$ respectively. It is assumed that the compound system’s evolution is driven by the Schrödinger equation, so the evolution operator is unitary.
Formally, an indirect measurement model for an observable $A$, introduced in [3] as a “measuring process”, is a quadruple
$({\cal K},|\xi\rangle,U,M),$ (1)
where $|\xi\rangle\in{\cal K}$ represents the apparatus state.
We explore the Heisenberg picture. To describe meter’s evolution, we represent it in the state space of the compound system, i.e., as $I\otimes M.$ The meter observable evolves as
$M(t)=U^{\star}(t)(I\otimes M)U(t).$ (2)
By the Born rule
$P(M(t)=x|\psi\xi)=\langle\psi\xi|E_{M(t)}(x)|\psi\xi\rangle.$ (3)
This is the probability distribution for the outputs of measurements done by the apparatus and given by the meter.
Definition. A measurement process $({\cal K},|\xi\rangle,U,M)$ reproduces the probability distribution for quantum observable $A$ (accurate von Neumann observable) if
$P(A=x|\psi)=P(M(T)=x|\psi\xi).$ (4)
Following [1], consider two remote observers $O_{1}$ and $O_{2}$ who perform joint measurements on a system ${\cal S},$ in mathematical terms it means that the meter quantum observables of the corresponding measurement processes commute,
$[M_{1}(t),M_{2}(t)]=0.$
Here each apparatus has its own state space, i.e., ${\cal K}={\cal K}_{1}\otimes{\cal K}_{2}.$ We call such measurements local. In this situation the joint probability distribution is well defined
$P(M_{1}(t)=x,M_{2}(t)=y|\psi\xi_{1}\xi_{2})=\langle\psi\xi_{1}\xi_{2}|E_{M_{1}(t)}(x)E_{M_{2}(t)}(y)|\psi\xi_{1}\xi_{2}\rangle.$ (5)
Theorem. (OIT [1]) Two observers performing the joint local and probability reproducible measurements of the same quantum observable $A$ on the system ${\cal S}$ should get the same outcome with probability 1:
$P(M_{1}(T)=x,M_{2}(T)=y|\psi\xi_{1}\xi_{2})=\delta(x-y)P(E=x|\psi)=\delta(x-y)\|E(x)\psi\|^{2}.$ (6)
4 Postulate on internally consistent description in the view of Ozawa Intersubjectivity Theorem
In OIT, as in postulate 6, observer $O_{1}$ measures $S,$ and observer $O_{2}$ also measures $S$ in the same basis. In our terms the later means that they measure the same observable $A.$ Then $O_{2}$ can interact with $O_{1}$ to “check the reading” of its pointer variable, the two values found are in agreement.
The important condition which is missed in postulate 6 is the probability reproducibility condition, only under this conditions outcomes of measurements performed by $O_{1}$ and $O_{2}$ coincide. Therefore, it is natural to complete postulate 6 by this condition. (But this is just my viewpoint and may be Carlo Rovelli would not support it.)
Moreover, due to OIT observer $O_{2}$ even need not to perform a measurement on $O_{1}.$ If observers are sure in validity of quantum theory, then they can be sure that they get the same outcome.
I suggest the following modification of postulate 6:
Postulate 6[OIT]. Under the condition of probability reproducibility, in a scenario where $O_{1}$ measures $S,$ and $O_{2}$ also measures $S$ in the same basis, the two values found are in agreement.
Acknowledgments
This study was partially supported by the EU-project: CA21169 - Information, Coding, and Biological Function: the Dynamics of Life (DYNALIFE). I would like to thank Masanao Ozawa for numerous lessons on quantum measurement theory, discussions on intersubjectivity of quantum measurements, and hospitality during my visits to Nagoya during the last 20 years.